Fri 26 Feb - Some issues with the 'lancaster' WiFi Network, and Fake Exchange Settings Message
'lancaster' WiFi Network
Windows users attempting to log-in to the 'lancaster' are not being re-directed to the login page and therefore cannot use the service. The Eduroam WiFi network is unaffected and so users could switch to using Eduroam. The Eduroam WiFi
Fake Exchange Settings Message
A fake message about Exchange settings, subject: "A new settings file for the (varies)@exchange.lancs.ac.uk has just be released" is circulating. Do NOT open the attachment. . More Information
Thu 25 Feb - All systems ran without problems
Wed 24 Feb - WebMail/UnixMail Maintenance
Webmail/Unixmail will be unavailable between 14:00 and 16:40 on Wednesday 24th February 2010 for maintenance. All the mailboxes are being migrated from one system to another as part of the ISS Decant. No user action is required. The finish time was extended to 17:30.
Tue 23 Feb - Telephone maintenance in Biology, LEC, Management School & Library
07:00-08:45: During the 'network at risk' period on Tuesday 23rd February ISS performed routine maintenance to telephony equipment serving the Biology, LEC, Management School & Library buildings. There was a short outage of up to 5 minutes to all telephone functionality from extensions beginning with a 1.
Mon 22 Feb - All systems ran without problems
Fri 19 Feb - All systems ran without problems
Thu 18 Feb - All systems ran without problems
Wed 17 Feb - Exchange Email Service Maintenance
The Exchange email system was temporarily unavailable, for the application of security updates, between 13:00 and 16:00 on Wednesday 17th February. Individuals saw interruption of service for no more than 30 minutes during this period.
Tue 16 Feb - Central Filestore Emergency Maintenance
Following advice from the vendor, ISS carried out emergency maintenance on the central filestore between 19:30 and 20:15 to prevent a more serious interruption to the service.
Mon 15 Feb - Central Filestore Problem
ISS investigated a problem with access to central filestore at around 20:00. Access to central filestore was restored at 20:35pm